Please show the formulas and steps. I am using office 360o o T А B C С D E G H 1 J K L M N Р R S 1 Name Score 1 Score 2 - Total Score - Pass/Fail Column1 Columi - 2 Den 28 32 3 Scott
Answer #1
Name Score 1 Score 2 Total Score Pass/Fail Column 1 Column 2
Den 28 32 60 Pass Poor You performed Fantastic
Scott 19 29 48 Fail Satisfactory You performed okay
Colin 13 33 46 Fail Satisfactory You performed okay
Dustin 27 34 61 Pass Good You performed Fantastic
Mike 20 30 50 Pass Satisfactory You performed okay
James 25 50 75 Pass Good You performed Fantastic
Yvette 15 48 63 Pass Good You performed Fantastic
Karl 27 39 66 Pass Good You performed Fantastic
Frank 14 28 42 Fail Satisfactory You performed okay

Assume above data in excel and below are the formulae's for reference:-

Formulae for 1) - B2+C2

2) =IF(OR(AND(B2>=20,C2>=25), AND(B2>=15,C2>=30)),"Pass","Fail")

3) =IF(B2+C2>60,"Good", IF(AND(B2+C2>40,B2+C2<60),"Satisfactory","Poor"))

4) =SWITCH(TRUE, D2>=60,"You performed Fantastic","You performed okay")

Also attached the screen clippping for your reference:

E2 =IF(OR(AND(B2>=20,C2>=25), AND(B2>=15,C2>=30)),Pass,Fail) H OH А 1 Name 2 Den 3 Scott 4 Colin 5 Dustin 6 Mike 7 James

F2 fx =IF(B2+C2>60,Good, IF(AND(B2+C2>40,B2+C2<60), Satisfactory, Poor)) H Formula B: B с D E F G 1 Name Score 1 Score